OpenStudy (anonymous):

adults tickets for the senior play cost$5 each and student tickets cost $3. A total of 1060 tickets sold costing $4500 were sold. How many student tickets were sold?

OpenStudy (nurali):

a= number of adult tickets t= number of student tickets 5a+3t=4500 a+t=1060 The substitution method: t=1060-a 5a+3(1060-a)=4500 5a+3180-3a=4500 2a=1320 a=660 t=400

OpenStudy (nurali):

Welcome to Openstudy